A company that manufactures magnetic membrane switches is investigating three production options that have the estimated cash flows below in millions of dollars if the interest rate of 15% per year, what is the annual worth of the In-house option in millions of dollars?

In-house

License

Contract

First cost, $

-30

-3

0

Annual operating cost, $ per year

-5

-0.2

-2

Annual income, $ per year

14

1.5

2.5

Salvage value, $

7

-

-

Life, years

10

5
Answer #1

Option : In-house

Initial Cost=Co=-$30 M

Annual operating cost per year=OM=-$5 M

Annual income per year=I=$14 M

Net Income per year=R=(OM+I)=(-5+14)=$9 M

Salvage Value=S=7 years

Useful life=n=10 years

AW=Co*(A/P,15%,10)+R+S*(P/F,15%,10)*(A/P,15%,10)

(P/F,15%,10)=1/(1+15%)^10=0.247185

(A/P, i, n) = 1-

0.15 (A/P, 0.15, 10) = 1-TTAH_ = 0.199252 1+0.15)10

AW=-30*0.199252+9+7*0.247185*0.199252 =$3.367205 M